# Automated Model Tuning for Multifidelity Uncertainty Propagation in Trajectory Simulation

### v1

@@ 1 Introduction
Multifidelity uncertainty propagation has been shown to enable more precise and efficient estimates of a trajectory’s quantities of interest (QoIs) [3] when compared to the standard Monte Carlo (MC) approach. Multifidelity uncertainty propagation techniques leverage low-fidelity models (fast but biased) to deliver efficient statistical estimates of a high-fidelity model’s (slow but accurate) QoIs. Multifidelity uncertainty propagation started, in the sampling context, with Heinrich [4] and Giles [5] who pioneered the use of coarser approximations for the solution of stochastic integral equations and stochastic differential equations, respectively. Building upon these seminal contributions, a large multi-level MC (MLMC) literature developed, see Giles’ review paper [6]. Several other related approaches were later proposed over the last decade, including multi-index MC [7, 8], multi-fidelity MC (MFMC) [9, 10, 11], multi-level control variate MC approaches [12, 13, 14] and, most recently, Approximate Control Variate (ACV) [15, 16, 17] approaches and multi-level best linear unbiased estimators (MLBLUE) [18, 19]. Despite differences in the construction of the classes of multi-fidelity estimators described above, virtually all of them rely on low-fidelity models directly derived from the high-fidelity model. For example, it is customary to use less accurate time integration, grid coarsening, and/or introduction of simplified physics. As such, low-fidelity models typically have hyperparameters that can affect both their accuracy and computational cost. The selection of these hyperparameters play an important part in efficiency and precision of multifidelity estimates [20, 21]. Low-fidelity model tuning (optimal selection of hyperparameters) is thus an important step to multifidelity uncertainty propagation [^1] An alternative approach to achieving similar goals is to rely on the construction of shared low-dimensional spaces to augment the correlation among low- and high-fidelity models through a mapping of the input spaces, see  [22, 23, 24, 25] for more details..
Though automated low-fidelity model tuning has recently been developed [20, 21] for trajectory uncertainty propagation, a number of practical questions still remain. For instance, in cases such as the use of onboard trajectory optimization (considering uncertainty) the computational cost of the model tuning must be incorporated in the available computational budget. In other words, one must choose to spend computation tuning the low-fidelity model(s) or to proceed with estimation using a sub-optimal low-fidelity model. The current work aims to elucidate this point to offer a practical guide to the application of model tuning.
@@
The Adaptable, Deployable Entry and Placement Technology (ADEPT) project test flight [26, 27] is used as a case study to develop practical recommendations for multifidelity model tuning. This case study uses the Program to Optimize Simulated Trajectories II (POST2), a legacy NASA Langley Research Center code, as a variable-fidelity simulator [28]. The goal of the case study is to estimate QoIs such as landing location and touchdown time given a set of uncertainties associated with initial vehicle state, atmospheric conditions, etc., under a limited computational budget. To simulate the use of onboard trajectory simulation (where estimates must be achieved on regular intervals of the flight), the budget includes any computation spent for the tuning of models. The development and use of Sandia National Laboratory’s flagship uncertainty quantification code Dakota [29] is the basis for the model tuning portion of this work.
2 The Multifidelity Model Tuning Problem
Control variate multifidelity MC uncertainty propagation centers on the approximation of the expected value (or any other moment) of a QoI, $Q$, of a (high-fidelity) system given access to $M$ low-fidelity approximations of $Q$: $Q_{i}$ for $i\in{1,2,..,M}$. As described below, the key problem in multifidelity modeling is to optimally allocate computational resources across the available models, i.e., how to determine the number of samples to evaluate for each model such that the estimator variance is minimized. In this work, this framework is expanded by considering that each low-fidelity approximation (model) $Q_{i}$ can be a function of hyperparameters $\beta_{i}$ where $\underline{\beta}$ denotes the collection of all model hyperparameters: $\underline{\beta}={\beta_{1}\cup\beta_{2}\cup\dots\cup\beta_{M}}$. The high-fidelity model is assumed to be fixed and not tunable. The expectation of $Q$ is approximated with an estimator $\tilde{Q}$, i.e., $\mathbb{E}\left[Q\right]\approx\tilde{Q}(\underline{\beta})$. Assuming an unbiased estimator (as will be the case in this work), minimizing the mean-squared error of the estimator is equivalent to minimizing its variance, $\mathbb{V}ar\!\left[\tilde{Q}(\underline{\beta})\right]$. The model tuning problem is usually constrained so that the computational cost of the estimator, $\mathcal{C}$, is less than $C_{budget}$, a prescribed computational budget [^2] An alternate formulation to minimize computational cost given an accuracy constraint on $\mathbb{V}ar\!\left[\tilde{Q}(\underline{\beta})\right]$ is also common.. Thus, the model tuning problem is defined at a high level by
$\beta^{*}=\operatorname*{arg\,min}_{\beta}\mathbb{V}ar\!\left[\tilde{Q}(\underline{\beta})\right]\quad\text{s.t.}\quad\mathcal{C}(\underline{\beta})<C_{budget}$ (1)
As discussed in the previous section, there are several available choices for the multifidelity estimator $\tilde{Q}(\underline{\beta})$ [e.g., 15, 18, 30]. This work focuses on the application of the approximate control variates (ACV) framework [15] to the model tuning problem, though many of the results and insights presented here are independent of the choice of $\tilde{Q}(\underline{\beta})$. The ACV framework is pursued because of its flexibility: ACV has unified existing methods like MFMC [10] and MLMC [6], it encompasses a large group of possible estimators [16, 17], and it also has direct relationships with MLBLUE [18, 19].
In the ACV framework the multifidelity estimator is defined as
@@ 2 The Multifidelity Model Tuning Problem
$\begin{split}\tilde{Q}\left(\underline{\beta},\underline{\alpha},\underline{\mathbf{z}}\right)&=\hat{Q}(\mathbf{z})+\sum_{i=1}^{M}\alpha_{i}\left(\hat{Q}_{i}(\beta_{i},\mathbf{z}_{i}^{1})-\hat{Q}_{i}(\beta_{i},\mathbf{z}_{i}^{2})\right)\end{split}$ (2)
where $\hat{Q}$ represents a standard MC estimator based on input samples which are realizations of the vector-valued random input variable $Z\in\mathbb{R}^{d}$, e.g.,
$\hat{Q}(\mathbf{z})=\frac{1}{N}\sum_{z\in\mathbf{z}}Q(z).$ (3)
Here, $N$ is the cardinality of $\mathbf{z}$. The ACV estimator in Equation (2) is based on the difference of two MC estimators for each low fidelity model $i$ using samples $\mathbf{z}_{i}^{1}$ and $\mathbf{z}_{i}^{2}$ multiplied by a control variate weight $\alpha_{i}$. The total samples associated with low-fidelity model $i$ is denoted $\mathbf{z}_{i}=\mathbf{z}_{i}^{1}\cup\mathbf{z}_{i}^{2}$. The ACV formulation can be used to obtain several multifidelity estimators by prescribing different partitioning schemes for $\mathbf{z}_{i}$; several partitioning schemes are reported in Table 1.
@@
The total cost of the ACV estimator is described by
$\mathcal{C}(\underline{\beta},\underline{\mathbf{z}})=Nw+\sum_{i=1}^{M}N_{i}w_{i}(\beta_{i})$ (4)
where $w$ and $w_{i}$ are the cost of the high-fidelity model and low-fidelity model $i$, respectively. $N_{i}$ denotes the cardinality of $\mathbf{z}_{i}$. It is noted that in the context of model tuning, it is assumed that the high-fidelity model is fixed and the cost of the low-fidelity models is dependent on the hyperparameters, $w_{i}(\beta_{i})$. Furthermore, while Equation 4 assumes that all low-fidelity models are tunable, in general there may be two groupings of $M^{f}$ fixed and $M^{t}$ tunable models.
By defining notation $\Delta_{i}(\beta_{i},\mathbf{z}_{i}^{×})=\hat{Q}_{i}(\beta_{i},\mathbf{z}_{i}^{1})-\hat{Q}_{i}(\beta_{i},\mathbf{z}_{i}^{2})$, $\underline{\mathbf{z}}=\left(\mathbf{z},\mathbf{z}_{1},\dots,\mathbf{z}_{M}\right)$, and $\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}})=\left[\Delta_{1}(\beta_{i},\mathbf{z}_{1}^{×}),\dots,\Delta_{M}(\beta_{i},\mathbf{z}_{M})\right]^{\mathrm{T}}$ Equation (2) can be simplified to:
@@
It can be shown that the optimal values for control variate weights are [15]:
$\underline{\alpha}^{opt}=-\mathbb{C}ov\left[\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}}),\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}})\right]^{-1}\mathbb{C}ov\left[\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}}),\hat{Q}(\underline{\mathbf{z}})\right].$ (6)
Using the optimal weights, the variance of the corresponding ACV estimator is given by [21]:
$\begin{split}\mathbb{V}ar\!\left[\tilde{Q}\left(\underline{\beta},\underline{\mathbf{z}}\right)\right]&=\mathbb{V}ar\!\left[\hat{Q}(\underline{\mathbf{z}})\right]-\mathbb{C}ov\left[\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}}),\hat{Q}(\underline{\mathbf{z}})\right]^{\mathrm{T}}\mathbb{C}ov\left[\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}}),\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}})\right]^{-1}\mathbb{C}ov\left[\underline{\Delta}(\underline{\beta},\underline{\mathbf{z}}),\hat{Q}(\underline{\mathbf{z}})\right]\\ &=\mathbb{V}ar\!\left[Q\right]\left(\frac{1}{N}-\frac{1}{N}\left[\mathbf{f}(\underline{\mathbf{z}})\circ\underline{\rho}(\underline{\beta})\right]^{\mathrm{T}}\left[\mathbf{F}(\underline{\mathbf{z}})\circ\mathbf{P}(\underline{\beta})\right]^{-1}\left[\mathbf{f}(\underline{\mathbf{z}})\circ\underline{\rho}(\underline{\beta})\right]\right)\end{split}$ (7)
where the matrix $\mathbf{P}$ represents the correlations among low-fidelity models $P_{ij}=\mathbb{C}orr\left[Q_{i},Q_{j}\right]$ and the $\underline{\rho}$ vector represents correlations with the high-fidelity model $\underline{\rho}=\left[\mathbb{C}orr\left[Q,Q_{1}\right],\dots,\mathbb{C}orr\left[Q,Q_{M}\right]\right]^{\mathrm{T}}$. The $\mathbf{f}$ vector and $\mathbf{F}$ matrix are defined using the cardinality of the sets in $\underline{\mathbf{z}}$ as well as the cardinality of the intersections of those sets. Definitions of $\mathbf{f}$ vector and $\mathbf{F}$ as well as the complete derivation of Equation (7) can be found in the Appendix. For a fixed setting of low-fidelity hyperparameters, $\hat{\underline{\beta}}$, the standard ACV sample allocation optimization problem is given by:
$\underline{\mathbf{z}}^{*}=\operatorname*{arg\,min}_{\underline{\mathbf{z}}}\mathbb{V}ar\!\left[\tilde{Q}\left(\hat{\underline{\beta}},\underline{\mathbf{z}}\right)\right]\quad\text{s.t.}\quad\mathcal{C}(\hat{\underline{\beta}},\underline{\mathbf{z}})<C_{budget}.\\$ (8)
@@
A general solution of this standard ACV sample allocation problem is currently unavailable due to the difficult nature of the optimization. Current approaches find an optimum on restricted subdomains of $\underline{\mathbf{z}}$ (such as those in Table 1). The most general of these approaches, parametric ACVs [16], considers many such restricted subdomains through parametric construction. The construction technique is based on developing trees (hierarchies) describing the relations of all sets in $\underline{\mathbf{z}}$ and then determining the recursive relation that minimizes estimator variance. Though subdomains generated in this way do not completely cover the domain of all possible $\underline{\mathbf{z}}$, the increased coverage compared to the methods in Table 1 typically results in lower variance. This expanded search for optimal sample allocation with parametric ACVs can be readily combined with automated model tuning, as demonstrated later.

### v2

@@ 1 Introduction
Multifidelity uncertainty propagation has been shown to enable more precise and efficient estimates of a trajectory’s quantities of interest (QoIs) [3] when compared to the standard Monte Carlo (MC) approach. Multifidelity uncertainty propagation techniques leverage low-fidelity models (fast but biased) to deliver efficient statistical estimates of a high-fidelity model’s (slow but accurate) QoIs. Multifidelity uncertainty propagation started, in the sampling context, with Heinrich [4] and Giles [5] who pioneered the use of coarser approximations for the solution of stochastic integral equations and stochastic differential equations, respectively. Building upon these seminal contributions, a large multi-level MC (MLMC) literature developed, see Giles’ review paper [6]. Several other related approaches were later proposed over the last decade, including multi-index MC [7, 8], multifidelity MC (MFMC) [9, 10, 11], multi-level control variate MC approaches [12, 13, 14] and, most recently, Approximate Control Variate (ACV) [15, 16, 17] approaches and multi-level best linear unbiased estimators (MLBLUE) [18, 19]. Despite differences in the construction of the classes of multifidelity estimators described above, virtually all of them rely on low-fidelity models directly derived from the high-fidelity model. For example, it is customary to use less accurate time integration, grid coarsening, and/or introduction of simplified physics. As such, low-fidelity models typically have hyperparameters that can affect both their accuracy and computational cost. The selection of these hyperparameters directly impacts the efficiency and precision of multifidelity estimators, making low-fidelity model tuning (i.e., the automated selection of optimal hyperparameters) an important avenue of research [^1] An alternative (and possibly complementary) approach to achieving similar goals is to rely on the construction of shared low-dimensional spaces to augment the correlation among low- and high-fidelity models through a mapping of the input spaces, see  [20, 21, 22, 23, 24, 25] for more details..
Automated model tuning has been recently demonstrated in the context of trajectory simulation [26, 27], thermal battery analysis [28], and high-speed flows [29] but the area remains largely unexplored beyond these studies. Bomarito et al. [26] first posed model tuning as an optimization problem where model hyperparameters that minimize multifidelity estimator variance are sought. However, this work mainly focused on exploring the potential increase in speed and precision for multifidelity trajectory simulation rather than providing a practical implementation of the approach. Thompson et al. [27] and Adams et al. [28] later employed Efficient Global Optimization (EGO) as a robust solution strategy, but both studies considered the model tuning optimization problem in isolation rather than analyzing its impact on cost and precision within a budget-constrained multifidelity uncertainty propagation analysis. In particular, multifidelity estimates of statistics were not actually formed using the tuned low-fidelity models; instead the performance was gauged through the optimal solutions for projected estimator variance. Furthermore, the trajectory simulation study of Thompson et al. [27] was limited to single tunable hyperparameter and only considered one QoI. Thus, a comprehensive study of the cost-versus-precision tradeoff provided by allocating computational budget to automated model tuning is lacking and its potential for improving multifidelity trajectory simulation is not fully understood.
@@
This paper provides an in-depth study of automated model tuning for multifidelity uncertainty propagation, representing the culmination of a line of previous work on the topic [26, 27]. A real-world entry, descent, and landing example - the Adaptable, Deployable Entry and Placement Technology (ADEPT) project test flight - [30, 31] is considered to develop practical recommendations for the approach. This case study uses the Program to Optimize Simulated Trajectories II (POST2), a legacy NASA Langley Research Center code, as a variable-fidelity simulator [32]. The goal here is to estimate QoIs such as landing location and touchdown time given a set of uncertainties associated with initial vehicle state, atmospheric conditions, etc., under a limited computational budget where upfront tuning costs cannot be ignored. After optimal hyperparameters are identified with model tuning, QoI estimators are formed using an iterated ACV approach [28] that adaptively estimates model QoI statistics (correlations, variances) that are required for multifidelity estimation. To explore the cost-versus-precision tradeoff associated with model tuning, results are benchmarked against baseline ACV estimators that use hand-tuned models [3] as well as the best-case ACV performance where optimal model hyperparameters and QoI statistics are known a priori. Sandia National Laboratories’ flagship uncertainty quantification code, Dakota [33], is used as the basis for the implementation of both ACV estimators and model tuning optimization. The primary contributions of this work are summarized as follows:
  • A rigorous study of the cost-versus-precision tradeoff from allocating upfront computation to automated model tuning by comparing the approach to both baseline and best-case performance within an end-to-end uncertainty propagation analysis for varying computational budgets.
  • An improvement to the state of the art for multifidelity trajectory simulation estimators demonstrated through comparisons to original work in this area [3].
  • The first demonstration of an online multifidelity uncertainty propagation procedure where neither optimal low-fidelity model hyperparameters nor knowledge of model statistics are known a priori using automated model tuning and iterated ACV.
  • Recommendations for applying automated model tuning to practical problems and avenues for future work to improve performance.
2 Multifidelity Estimators and Model Tuning
In this section, the Approximate Control Variate (ACV) multifidelity framework is briefly introduced as based on previous work [15, 16]. Next, the so-called model tuning problem, which constitutes the main contribution of the paper, is introduced and discussed. It is important to note that the model tuning idea can be adopted with minor modification in virtually any existing multifidelity sampling estimator, e.g., MFMC [10] or MLBLUE [18, 19], but the ACV framework is emphasized here due to its flexibility. ACV has unified existing methods, MLMC [6] and MFMC, it encompasses a large group of possible estimators [16, 17], and it also has direct relationships with MLBLUE [18, 19, 17]. The idea of all multifidelity estimators is to optimally combine realizations from an ensemble of models with varying cost and accuracy to increase the statistical precision of targeted quantities for the high-fidelity model, for which only a limited number of simulations can be acquired. In this presentation, the focus is on the estimation of an expected value, but the extension to higher order moments or densities is also possible [34, 35, 36].
Let $Q:\mathbb{R}^{d}\rightarrow\mathbb{R}$ represent the mapping from a vector of input $Z\in\mathbb{R}^{d}$ to a scalar quantity of interest (QoI) for a high-fidelity model, while $Q_{i}$ represent $M$ suitable low-fidelity approximations of $Q$. An Optimal Control Variate (OCV) estimator  [37, 38, 39], is a Monte Carlo-based linear control variate estimator for the expected value of $Q$, hereafter denoted as $\mathbb{E}\left[Q\right]$, with respect to $Z$. It can be written as
$\hat{Q}^{OCV}=\hat{Q}(\mathbf{z})+\sum_{i=1}^{M}\alpha_{i}\left(\hat{Q}_{i}(\mathbf{z})-\mu_{i}\right),$ (1)
where the symbol $\hat{\cdot}$ denotes a MC sampling estimator, e.g.,
$\hat{Q}(\mathbf{z})=\frac{1}{N}\sum_{z\in\mathbf{z}}Q(z)$ (2)
which is based on a set, $\mathbf{z}$, of $N$ independent realizations of $Z$ (identical definitions apply for $Q_{i}$ with $i=1,\dots,M$), $\alpha_{i}$ represent weights to be optimized, and $\mu_{i}$ are the known expected values for the $M$ low-fidelity models, i.e., $\mathbb{E}\left[Q_{i}\right]=\mu_{i}$. Sampling estimators are random variables because they are based on finite realizations of the input, but it is easy to show that they are unbiased, i.e., $\mathbb{E}\left[\hat{Q}^{OCV}\right]=\mathbb{E}\left[\hat{Q}(\mathbf{z})\right]=\mathbb{E}\left[Q\right]$ for any choice of $\alpha_{i}\in\mathbb{R}$. On the other hand, the OCV variance can be minimized when the weights are chosen as $\underline{\alpha}=\left[\alpha_{1},\dots,\alpha_{M}\right]^{\mathrm{T}}=-\mathbf{C}^{-1}\mathbf{c}$, where $\mathbf{C}\in\mathbb{R}^{M\times M}$ denotes the covariance matrix among $Q_{i}$ and $\mathbf{c}\in\mathbb{R}^{M}$ denotes the covariance vector between $Q$ and $Q_{i}$. This variance can be expressed as
$\mathbb{V}ar\!\left[\hat{Q}^{OCV}\right]=\mathbb{V}ar\!\left[\hat{Q}(\mathbf{z})\right]\left(1-\frac{\mathbf{c}^{\mathrm{T}}\mathbf{C}^{-1}\mathbf{c}}{\mathbb{V}ar\!\left[Q\right]}\right)=\mathbb{V}ar\!\left[\hat{Q}(\mathbf{z})\right]\left(1-R^{2}_{OCV}\right),$ (3)
where $\mathbb{V}ar\!\left[\hat{Q}(\mathbf{z})\right]=\dfrac{\mathbb{V}ar\!\left[Q\right]}{N}$ is the MC estimator variance and $0\leq R^{2}_{OCV}\leq 1$ is a measure of the variance reduction with respect to the MC estimator’s variance. The OCV framework demonstrates that, by providing correlated low-fidelity models, it is possible in principle to reduce the MC variance; however, for realistic problems, the expected values $\mu_{i}$ need to be estimated and a sampling strategy needs to be chosen.
The ACV framework encompasses several estimators and is defined by using $N_{i}=r_{i}N$ samples for each low-fidelity model and by partitioning this set in two ordered subsets $\mathbf{z}_{i}^{1}$ and $\mathbf{z}_{i}^{2}$ so that $\mathbf{z}_{i}=\mathbf{z}_{i}^{1}\cup\mathbf{z}_{i}^{2}$. The subsets are not required to be disjoint, i.e., it is admissible to have $\mathbf{z}_{i}^{1}\cap\mathbf{z}_{i}^{2}\neq 0$. The generic ACV estimator can then be written as
@@ 2 Multifidelity Estimators and Model Tuning
$\begin{split}\tilde{Q}\left(\underline{\alpha},\underline{\mathbf{z}}\right)&=\hat{Q}(\mathbf{z})+\sum_{i=1}^{M}\alpha_{i}\left(\hat{Q}_{i}(\mathbf{z}_{i}^{1})-\hat{Q}_{i}(\mathbf{z}_{i}^{2})\right)\\ &=\hat{Q}(\mathbf{z})+\underline{\alpha}^{\mathrm{T}}\underline{\Delta},\end{split}$ (4)
where the vector of difference estimators is defined as $\underline{\Delta}=\left[\Delta_{1}=\hat{Q}_{i}(\mathbf{z}_{1}^{1})-\hat{Q}_{i}(\mathbf{z}_{1}^{2}),\dots,\Delta_{M}=\hat{Q}_{i}(\mathbf{z}_{M}^{1})-\hat{Q}_{i}(\mathbf{z}_{M}^{2})\right]^{\mathrm{T}}$ and each term is obtained as the difference of two MC estimators for the $i$th model based on $\mathbf{z}_{i}^{1}$ and $\mathbf{z}_{i}^{2}$. Several estimators can be obtained by prescribing pre-determined relationships between $\underline{\mathbf{z}}$, $\mathbf{z}_{i}^{1}$, and $\mathbf{z}_{i}^{2}$; MLMC, MFMC, and two ACV variants, namely ACV-IS and ACV-MF, are defined in Table 1 (from [15]). It is important to note that MLMC is recovered by introducing the additional simplification of fixing the weights $\underline{\alpha}=-1$, whereas a weighted MLMC estimator can be obtained if these weights are optimized while still keeping the same MLMC partitioning for $\mathbf{z}_{i}^{1}$, and $\mathbf{z}_{i}^{2}$, as discussed in [15].
@@
It can be shown that the optimal values for control variate weights are [15]
$\underline{\alpha}^{opt}=-\mathbb{C}ov\left[\underline{\Delta}(\underline{\mathbf{z}}),\underline{\Delta}(\underline{\mathbf{z}})\right]^{-1}\mathbb{C}ov\left[\underline{\Delta}(\underline{\mathbf{z}}),\hat{Q}(\underline{\mathbf{z}})\right],$ (5)
while the corresponding ACV variance is given by [27]
$\begin{split}\mathbb{V}ar\!\left[\tilde{Q}\left(\underline{\mathbf{z}}\right)\right]&=\mathbb{V}ar\!\left[\hat{Q}(\underline{\mathbf{z}})\right]-\mathbb{C}ov\left[\underline{\Delta}(\underline{\mathbf{z}}),\hat{Q}(\underline{\mathbf{z}})\right]^{\mathrm{T}}\mathbb{C}ov\left[\underline{\Delta}(\underline{\mathbf{z}}),\underline{\Delta}(\underline{\mathbf{z}})\right]^{-1}\mathbb{C}ov\left[\underline{\Delta}(\underline{\mathbf{z}}),\hat{Q}(\underline{\mathbf{z}})\right]\\ &=\frac{\mathbb{V}ar\!\left[Q\right]}{N}\left(1-\left[\mathbf{f}(\underline{\mathbf{z}})\circ\underline{\rho}\right]^{\mathrm{T}}\left[\mathbf{F}(\underline{\mathbf{z}})\circ\mathbf{P}\right]^{-1}\left[\mathbf{f}(\underline{\mathbf{z}})\circ\underline{\rho}\right]\right)\\ &=\frac{\mathbb{V}ar\!\left[Q\right]}{N}\left(1-R^{2}_{ACV}\right)\end{split}$ (6)
where $\mathbf{P}$ is the matrix of the correlations among low-fidelity models (i.e., $P_{ij}=\mathbb{C}orr\left[Q_{i},Q_{j}\right]$), the $\underline{\rho}$ vector represents correlations with the high-fidelity model (i.e., $\underline{\rho}=\left[\mathbb{C}orr\left[Q,Q_{1}\right],\dots,\mathbb{C}orr\left[Q,Q_{M}\right]\right]^{\mathrm{T}}$), and $\circ$ denotes the Hadamard (element-by-element) product. The $\mathbf{f}$ vector and $\mathbf{F}$ matrix are defined using the cardinality of the sets in $\underline{\mathbf{z}}$ as well as the cardinality of the intersections of those sets. Definitions of $\mathbf{f}$ and $\mathbf{F}$ as well as the complete derivation of Equation (6) can be found in the Appendix. The variance reduction of ACV is bounded by the variance reduction of OCV, i.e., $0\leq R^{2}_{ACV}<R^{2}_{OCV}\leq 1$, and it was demonstrated in [15] that the main advantage of ACV estimators like ACV-MF and ACV-IS (which are non recursive) is that for $r_{i}\rightarrow\infty$, $R^{2}_{ACV}\rightarrow R^{2}_{OCV}$, while recursive approaches like MLMC and MFMC are limited by the largest correlation within the low-fidelity model ensemble. Further details are provided in  [15].
2.1 Generalized ACV Estimators
While ACV estimators like ACV-MF and ACV-IS are optimal in the limiting case of $r_{i}\rightarrow\infty$, recursive estimators can still be advantageous for small $r_{i}$ regimes. Therefore, the optimal ACV estimator depends on the available data and generalized ACV estimators were introduced in [26]. These generalized ACV estimators allow for the restructuring of the control variates according to a graph, where the graphs are defined as a series of directed connections $i\rightarrow j$ that indicate that the model $Q_{i}$ acts as a control variate of $Q_{j}$. Recursive estimators like MLMC and MFMC correspond to an ordered sequence, while ACV estimators like ACV-MF and ACV-IS correspond to a peer structure; generalized ACV estimators can encode graphs that simultaneously include ordered sequences and peer structures. From a numerical standpoint, when the number of low-fidelity models is small, the best graph can be selected through enumeration, i.e., all solutions are computed and the one with the smallest variance is selected. The graph effect is represented by a different representation of the vector $\mathbf{f}$ and the matrix $\mathbf{F}$; see [26] for additional details. In the next section the resource allocation is presented for any ACV estimator, given a fixed set of models, a prescribed sampling scheme, and a specific graph.
2.2 The Resource Allocation Problem
The most important aspect of multifidelity sampling estimators is that computational resources can be optimally allocated across the ensemble of models to minimize the estimator variance for a prescribed computational budget [^2] An alternate formulation to minimize computational cost given an accuracy constraint on $\mathbb{V}ar\!\left[\tilde{Q}(\underline{\mathbf{z}})\right]$ is also common.. Let the total cost $\mathcal{C}$ of an ACV estimator
$\mathcal{C}(\underline{\mathbf{z}})=Nw+\sum_{i=1}^{M}N_{i}w_{i},$ (7)
where $w$ and $w_{i}$ are the cost of the high-fidelity model and low-fidelity model $i$, respectively. The resource allocation problem is formulated as
$\underline{\mathbf{z}}^{*}=\operatorname*{arg\,min}_{\underline{\mathbf{z}}}\mathbb{V}ar\!\left[\tilde{Q}\left(\hat{\underline{\mathbf{z}}}\right)\right]\quad\text{s.t.}\quad\mathcal{C}(\underline{\mathbf{z}})<C_{budget},\\$ (8)
@@ 2.2 The Resource Allocation Problem
A general solution of this standard ACV sample allocation problem is currently unavailable due to the difficult nature of the optimization. Current approaches find an optimum on restricted subdomains of $\underline{\mathbf{z}}$ (such as those in Table 1). The most general of these approaches, parametric ACVs [16], considers many such restricted subdomains through parametric construction. The construction technique is based on developing trees (hierarchies) describing the relations of all sets in $\underline{\mathbf{z}}$ and then determining the recursive relation that minimizes estimator variance. Though subdomains generated in this way do not completely cover the domain of all possible $\underline{\mathbf{z}}$, the increased coverage compared to the methods in Table 1 typically results in lower variance.
2.3 The Model Tuning Problem
In this section, the main contribution of the paper is introduced. As described above, the key problem in multifidelity modeling is to optimally allocate computational resources across the available models. In this work, this framework is expanded by considering that each low-fidelity approximation (model) $Q_{i}$ can be a function of hyperparameters $\beta_{i}$ where $\underline{\beta}$ denotes the collection of all model hyperparameters: $\underline{\beta}={\beta_{1}\cup\beta_{2}\cup\dots\cup\beta_{M}}$. The high-fidelity model is assumed to be fixed and not tunable in order to provide for an invariant reference model [^3] Accounting for additional accuracy/cost resulting from high-fidelity tuning would typically require an additional reference, e.g. experimental data, which introduces an additional information source and expands the scope of the ensemble.. Furthermore, for simplicity in the following, it is assumed that all low-fidelity models have hyperparameters, but this is not necessary and some low-fidelity models could be fixed. The expectation of $Q$ is approximated with an estimator $\tilde{Q}$, i.e., $\mathbb{E}\left[Q\right]\approx\tilde{Q}(\underline{\mathbf{z}},\underline{\beta})$, where now an explicit dependency on the hyperparameter vector is introduced. To capture the effect of hyperparameters on the costs of the low-fidelity models (e.g., $\underline{\beta}$ can include a control on the spatial resolution of a low-fidelity model), Eq. 7 is modified to include the dependency of hyperparameters,
$\mathcal{C}(\underline{\beta},\underline{\mathbf{z}})=Nw+\sum_{i=1}^{M}N_{i}w_{i}(\beta_{i}).$ (9)
The model tuning problem can now be formulated by augmenting the resource allocation problem (Eq. (8)) with a simultaneous search for the optimal hyperparameters [^4] See Sec. 2.5 for additional formulation details. as follows:
$\underline{\mathbf{z}}^{*},\beta^{*}=\operatorname*{arg\,min}_{\underline{\mathbf{z}},\beta}\mathbb{V}ar\!\left[\tilde{Q}(\underline{\mathbf{z}},\underline{\beta})\right]\quad\text{s.t.}\quad\mathcal{C}(\underline{\mathbf{z}},\underline{\beta})<C_{budget},$ (10)
where the ACV estimator is now also a function of the hyperparameters through the control variates,
[…]
and its variance follows directly from Eq. 6, i.e., $\mathbb{V}ar\!\left[\tilde{Q}\left(\underline{\mathbf{z}},\underline{\beta}\right)\right]=\mathbb{V}ar\!\left[\hat{Q}(\mathbf{z})\right]\left(1-R^{2}_{ACV}(\underline{\beta})\right)$ with
$R^{2}_{ACV}(\underline{\beta})=\left[\mathbf{f}(\underline{\mathbf{z}})\circ\underline{\rho}(\underline{\beta})\right]^{\mathrm{T}}\left[\mathbf{F}(\underline{\mathbf{z}})\circ\mathbf{P}(\underline{\beta})\right]^{-1}\left[\mathbf{f}(\underline{\mathbf{z}})\circ\underline{\rho}(\underline{\beta})\right].$ (12)
It should be noted that the optimal weights in Eq. 5 and the vector $\underline{\rho}$ and the matrix $\mathbf{P}$ in Eq. 12 depend on $\underline{\beta}$ since $\rho_{i}(\beta_{i})=\mathbb{C}orr\left[Q,Q_{i}(\beta_{i})\right]$ and $P_{ij}(\beta_{i},\beta_{j})=\mathbb{C}orr\left[Q_{i}(\beta_{i}),Q_{j}(\beta_{j})\right]$, respectively.